### v1

# Non-Hermitian topology of transport in Chern insulators

Abstract
It has recently been shown that signatures of non-Hermitian topology can be realized in a conventional quantum Hall device connected to multiple current sources. These signatures manifest as robust current-voltage characteristics, dictated by the presence of a nontrivial, non-Hermitian topological invariant of the conductance matrix. Chiral edge states are believed to be responsible for this non-Hermitian response, similar to how they lead to a quantized Hall conductivity in the presence of a single current source. Here, we go beyond this paradigm, showing that multi-terminal conductance matrices can exhibit non-Hermitian topological phase transitions that cannot be traced back to the presence and directionality of a boundary-localized chiral mode. By performing quantum transport simulations in the quantum Hall regime of monolayer graphene, we find that when the chemical potential is swept across the zeroth Landau level, unavoidable device imperfections cause the appearance of an additional non-Hermitian phase of the conductance matrix. This highlights graphene as an ideal platform for the study of non-Hermitian topological phase transitions, and is a first step towards exploring how the geometry of quantum devices can be harnessed to produce robust, topologically-protected transport characteristics.
Introduction — A physical system can have a non-Hermitian topological description if it is coupled to an environment and the resulting gains and losses are customized to reach a topological phase [1, 2]. Such a topological phase offers robust characteristics, with potential applications such as light funnels [3], amplifiers [4], and exponentially precise sensors [5, 6]. The experimental realization of these phases usually relies on platforms where gain and loss can be precisely controlled in each individual building block. Examples include ultracold atoms [7], optical systems [8, 9], and meta-materials [10, 11, 3, 12, 13, 14, 15].
A recent study takes an alternative approach, which leverages the quantum transport properties of a well-known Hermitian topological phase – the quantum Hall phase in a GaAs/AlGaAs heterostructure [16]. In the presence of multiple current sources, the current-voltage characteristics of the quantum Hall phase show robust features that originate from the non-Hermitian topology of the conductance matrix, not of the Hamiltonian. The chiral edge modes ensure that charge propagates between neighboring contacts only in one direction, meaning that the multi-terminal conductance matrix becomes numerically equivalent to the Hamiltonian matrix of the Hatano-Nelson (HN) chain [17], a prototypical model of non-Hermitian topology. In effect, each terminal plays the role of a ‘site’ in the HN chain, while the edge modes propagating between adjacent contacts play the role of nearest-neighbor, one-way ‘hoppings.’
Away from the quantum Hall plateau regime, the bulk becomes conducting and diffusive transport occurs between any pair of contacts, such that the analogy between chiral edge modes and nonreciprocal nearest-neighbor hoppings breaks down. All elements of the non-Hermitian conductance matrix become sizeable, potentially leading to additional topological phases and phase transitions that are not the direct consequence of the chiral modes. Here, we demonstrate the emergence of an additional phase in the non-Hermitian, multi-terminal conductance matrix of Chern insulators. This extra phase is not caused by boundary-localized chiral modes, arising instead in the transition region between two quantum Hall plateaus with opposite Chern numbers. The topologically trivial phase is due to the inhomogeneities present in the conductance matrix, which occur in any experimental setup. Our work shows that, beyond the quantum Hall regime, the quality and geometry of multi-terminal devices can play an important role in generating robust current-voltage characteristics protected by non-Hermitian topology.
To illustrate this concept, we turn to graphene, one of the most accessible and highly-tunable platforms for realizing quantum Hall physics. Using a toy model, we first highlight the correspondence between chiral edge modes and the non-Hermitian topology of the conductance matrix. This correspondence remains valid even as the chemical potential is swept through the charge neutrality point and the edge modes change direction. Next, we introduce inhomogeneities into the multi-terminal conductance matrix, showing that they cause the emergence of an additional phase.
Transport in graphene and its connection to non-Hermitian topology — We use a toy model for graphene consisting of spinless fermions hopping on a honeycomb lattice. The Hamiltonian takes the form
$H=t\sum_{\langle ij\rangle}^{{\phantom{{\dagger}}}}c^{{\dagger}}_{i}c^{{\phantom{{\dagger}}}}_{j}+\mu\sum_{i}c_{i}^{{\dagger}}c^{{\phantom{{\dagger}}}}_{i},$ (1)
[…]
We introduce a perpendicular magnetic field by changing the vertical hoppings in Fig. 1(a) from $t$ to $t\exp(i\phi n_{x}^{{\phantom{{\dagger}}}})$, where $\phi/2\pi$ is the flux per hexagon, in units of the flux quantum. This gaps out the Dirac cones and results in the formation of Landau levels, symmetrically distributed around the charge neutrality point, $E=0$ [18] (see the supplemental material (SM) for details [19]). Since the model is spinless, sweeping the chemical potential from $\mu>0$ to $\mu<0$ reverses the edge mode propagation direction from counter-clockwise to clockwise, leading to a direct transition between phases with Chern numbers $+1$ and $-1$.
Figure 1: Panel (a): Graphene lattice with two sites (shown as white and black circles) in each unit cell (marked by the yellow dashed contour). $\vec{e}_{x}^{{\phantom{{\dagger}}}}$ and $\vec{e}_{y}^{{\phantom{{\dagger}}}}$ are the Bravais vectors and $t$ is the nearest neighbor hopping. Panel (b): Schematic of a graphene disk with six contacts (blue) attached to its boundaries and labeled from one to six. The transmission between adjacent leads in the clockwise and anticlockwise directions is represented by red and green arrows, respectively. Panel (c): Conductance matrix $G$ for a disk of graphene with six leads [as in (b)], deep in a quantum Hall phase with Chern number $-1$ ($\mu/t=-0.2$, $\phi=0.18$). The conductance (in units of $e^{2}/h$) is quantized between adjacent leads in the downstream direction of the edge modes and is zero in the opposite direction. The full conductance matrix is equivalent to a Hatano-Nelson Hamiltonian under PBC whereas the sub-matrix (enclosed by blue dashed lines) is equivalent to the same under OBC. Panel (d): SPD plotted for the $G$ matrix in panel (c) against the site index (lead index) on a log-linear scale under OBC (non-Hermitian skin effect) and PBC.
Using the Kwant package [20], we attach $N$ equally-spaced ideal leads to the boundary of a finite-sized, disk-shaped system, as shown in Fig. 1(b) for $N=6$. Numerical details are described in the SM [19]. This models an experimental setup in which $N$ contacts are connected to a graphene sample. In the linear regime, the currents injected into each of the contacts, $\vec{I}=(I_{1},\ldots,I_{N})$, are related to the potentials on each of the contacts, $\vec{V}=(V_{1},\ldots,V_{N})$, as $\vec{I}=G\vec{V}$, where $G$ is the $N\times N$ conductance matrix. At zero-temperature and in the zero bias voltage limit, the elements of $G$ take the form
$G_{ij}=\frac{e^{2}}{h}\left[\delta_{ij}{\cal N}_{j}-{\rm tr}\left(S^{\phantom{{\dagger}}}_{ij}S^{\dagger}_{ij}\right)\right].$ (2)
Here, $e$ is the electric charge, $h$ is Planck’s constant, $\delta_{ij}$ denotes the Kronecker delta, ${\cal N}_{j}$ is the number of propagating modes in lead $j$, and ${\rm tr}$ denotes the trace. The matrix $S_{ij}$ is a block of the full, multi-terminal scattering matrix, containing the probability amplitudes for a particle to be transmitted from any incoming mode of lead $j$ to any outgoing mode of lead $i$.
For the six-terminal system shown in Fig. 1(b), setting $\mu=-0.2t$ and $\phi=0.18$, such that the system is deep in the quantum Hall phase with Chern number $-1$, leads to the conductance matrix shown in Fig. 1(c). This matrix is identical to the non-Hermitian Hamiltonian matrix of the Hatano-Nelson tight-binding model [17], a chain of orbitals connected by nonreciprocal nearest-neighbor hoppings. Thus, each of the diagonal elements $G_{ii}$ can be thought of as being analogous to the onsite energy of a site in the fictitious chain. The off-diagonal elements are analogous to hopping integrals connecting the different fictitious sites. In this case, each of the sites has an onsite energy equal to $+1$, one-way hoppings equal to $-1$ connect the $j^{\rm th}$ and $(j-1)^{\rm th}$ sites, and since $G_{61}=-1$, the chain has periodic boundary conditions (PBC). Realizing an analogue of the Hatano-Nelson model with open boundary conditions (OBC) can be done by grounding the $6^{\rm th}$ contact (setting the potential $V_{6}=0$). In this case the currents and voltages corresponding to the first five terminals would be related by the $5\times 5$ conductance sub-matrix, $G_{\rm OBC}$, shown by the dashed box in Fig. 1(c).
The nontrivial topology of the Hatano-Nelson Hamiltonian matrix, and thus of the multi-terminal conductance matrix, results in robust current-voltage characteristics. A transport regime in which the current vector is proportional to the voltage vector $\vec{I}\propto\vec{V}$ can only be reached provided that these vectors are eigenvectors of the conductance matrix. In the OBC case (grounded $6^{\rm th}$ contact), all eigenvectors are exponentially localized to one end of the fictitious Hatano-Nelson chain, a non-Hermitian topological phenomenon referred to as the non-Hermitian skin effect [21, 22]. This can be seen from the probability density summed over all right eigenvectors $v_{n}$ of the conductance matrix, ${\rm SPD}=\sum_{n}|v_{n}|^{2}$. Thus, when $\vec{I}\propto\vec{V}$ the currents and voltages on terminals $1$ to $5$ follow the exponentially decaying pattern shown in Fig. 1(d).
In contrast, for the PBC setup, when $\vec{I}\propto\vec{V}$ the six currents and voltages do not show an exponential profile. Both of these proportionality regimes can be reached in a robust and predictable way by means of iterative transport measurements, as shown experimentally in Ref. [16].
Inhomogeneity and topological phase transitions — We are interested in studying the topology of the conductance matrix away from the quantum Hall plateau regime, when bulk transport is significant and all elements of $G$ are nonzero. To this end, we follow Refs. [23, 16], and focus only on the OBC setup in the following. The $N^{\rm th}$ contact is grounded at all times and only the $(N-1)\times(N-1)$ sub-block of the conductance matrix, $G_{\rm OBC}$, is considered. We use a topological invariant, $w_{\rm PD}^{{\phantom{{\dagger}}}}$, that can be determined directly from this conductance matrix, and which predicts the existence of a non-Hermitian skin effect. To compute it, we first subtract the average of the diagonal elements from the conductance matrix, $\hat{G}=G_{\rm OBC}-\lambda\mathds{1}$, where $\lambda={\rm tr}(G_{\rm OBC})/(N-1)$ and $\mathds{1}$ is the identity matrix. Writing the polar decomposition of this matrix, $\hat{G}=QP$, where $Q$ is unitary and $P$ is a positive-definite matrix, yields
$w_{\rm PD}^{{\phantom{{\dagger}}}}(\hat{G})=\mathcal{T}(Q^{{\dagger}}[Q,X]),$ (3)
where $X={\rm diag}(1,2,\ldots,N-1)$ encodes the lead index, and $\mathcal{T}$ is the trace per unit volume evaluated over the middle indices, far from the ends (see SM for details [19]).
As expected, we confirm that $w_{\rm PD}^{{\phantom{{\dagger}}}}$ correctly predicts the appearance of a non-Hermitian skin effect in the conductance matrix of Fig. 1(d). Deep in the quantum Hall plateau regime, where the boundary-localized chiral modes produce one-way nearest-neighbor hoppings of the fictitious Hatano-Nelson chain, the invariant in Eq. (3) is locked to the Chern number. We have verified that sweeping the chemical potential across the charge neutrality point leads to a direct transition in the Chern number from $+1$ to $-1$, and simultaneously to a direct transition of $w_{\rm PD}^{{\phantom{{\dagger}}}}$, from $-1$ to $+1$. Correspondingly, there is a change of direction in the non-Hermitian skin effect: all eigenvectors of the conductance matrix are all localized either on the left or the right side of the chain, depending on the sign of $\mu$.
In any experiment, we expect that the conductance matrix will be inhomogeneous due to multiple factors. These include the varying quality of different contacts, unequal distances between them, as well as disorder and defects in the sample. For simplicity, we obtain inhomogeneous conductance matrices $G$ by randomly changing the potential on each site of the honeycomb lattice, $\mu\rightarrow\mu+\delta_{i}^{{\phantom{{\dagger}}}}$, where $\delta_{i}^{{\phantom{{\dagger}}}}$ is chosen from a uniform distribution $(-\delta,\delta)$.
The average non-Hermitian topological invariant of the conductance matrix can be expressed as:
$\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=\frac{\sum_{i=1}^{n}w_{\rm PD}^{{\phantom{{\dagger}}}}(\hat{G_{i}})}{n},$ (4)
for $n$ independent realizations of the random onsite potential, each producing a different conductance matrix $G_{i}$. Figure 2 shows that for $\delta=0$ the conductance matrix exhibits a direct transition from $w_{\rm PD}^{{\phantom{{\dagger}}}}=+1$ to $-1$ as the chemical potential is swept across the charge neutrality point, as mentioned above. Turning on inhomogeneities, however, leads to the appearance of a trivial phase of the conductance matrix, $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=0$, for which the current-voltage characteristics do not show the exponential profile of the non-Hermitian skin effect (see SM for details [19]). We observe that increasing $\delta$ leads to an enlargement of this phase, which appears for a wider range of chemical potential $\mu$.
Figure 2: $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle$ evaluated for conductance matrices of a monolayer spinless graphene disk with 10 leads, one of which is grounded (9 lead OBC matrix) as a function of varying $\delta/t$ and chemical potential $\mu/t$. The trivial phase ($\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=0$) becomes more prominent close to the charge neutrality point $\mu/t=0$ as $\delta$ increases. Here, $\phi=0.18$.
The intervening trivial non-Hermitian phase of the conductance matrix appears in a region where the bulk of the graphene is conducting, due to the fact that random potential fluctuations broaden the zeroth Landau level. However, it is not caused by the presence of conducting bulk modes. To show this, we consider a different averaging scheme, where we first average the conductance matrix over the random realizations of onsite disorder, and only afterwards compute the topological invariant:
$\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle_{2}^{{\phantom{{\dagger}}}}=w_{\rm PD}^{{\phantom{{\dagger}}}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{G_{i}}\right).$ (5)
This procedure leads to a conductance matrix with uniform entries on each diagonal, where the broadening of the zeroth Landau level is still present, but does not produce an extra non-Hermitian phase in the conductance matrix (see Fig. 3).
Figure 3: Panel (a): Absolute value of disorder-averaged conductances $\langle G\rangle_{1,2}$ (purple) and $\langle G\rangle_{2,1}$ (green), showing the transition between two quantum Hall phases with Chern numbers $C=1$ and $C=-1$, for a spinless graphene disk with $N=40$ contacts. In these phases, the conductance through the bulk is negligible, as indicated by the $\langle G\rangle_{1,20}\times 100$ (multiplied by a factor of 100 for ease of visibility) (blue). This bulk conductance becomes non-zero in the transition region between the two quantum Hall phases, indicating the diffusive transport regime [see SM for details [19]]. Panel (b): evaluated $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle$ shows the presence of a trivial phase close to $\mu/t=0$ between two non-Hermitian topological phases ($+1$ and $-1$). This is in contrast to $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle_{2}$ which shows a direct transition between the two phases. Here, $\delta/t=1.2$, and $\phi=0.18$. Panel (c): A $6\times 6$ conductance sub-matrix at $\mu=-0.04t$ for a single realization of onsite disorder, indicating the inhomogeneities in its entries. Panel (d): A $6\times 6$ sub-matrix of the disorder-averaged conductance matrix $\langle G\rangle$ at $\mu=-0.04t$ exhibiting the uniformity in its entries (see SM for details [19]).
The true reason for emergence of the extra phase between $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=\pm 1$ phases is due to inhomogeneities in the conductance matrix and is a natural consequence of the current conservation rules. Since the sum of all elements of $G$ over any column must vanish due to Kirchhoff’s laws, $\sum_{i=1}^{L}G_{ij}^{{\phantom{{\dagger}}}}=0$, random conductance fluctuations between different pairs of leads necessarily lead to random changes on the conductance matrix diagonal. On the level of the fictitious, 1D Hatano-Nelson chain simulated by $G$, this corresponds to the presence of onsite disorder, which is known to lead to an extra $\langle w_{\rm PD}^{{\phantom{{\dagger}}}}\rangle=0$ phase (see Ref. [23] and SM [19]).
Conclusion — We have theoretically investigated the non-Hermitian topological behavior associated with multiple-current-source transport in Chern insulators, using the quantum Hall phases of a honeycomb graphene model as an example. Our simulations demonstrated the presence and robustness of non-Hermitian phases in the conductance matrix of this system, which can easily be tuned through chemical potential variations. This variation can be achieved in practice via gate voltage adjustments, making graphene a promising platform to realize current-voltage relations protected non-Hermitian topology.
Interestingly, we found that inhomogeneities in the conductance matrix, which can arise, e.g. due to varying contact quality or inter-contact distance lead to the emergence of an additional, trivial non-Hermitian phase in the system’s conductance matrix. This implies that away from the quantum Hall plateau regime, the quality and geometry of quantum devices may play a decisive role in producing transport properties associated with non-Hermitian topology. Further, while we have focused on a graphene toy model for simplicity, we expect this behavior to be generic to a larger class of Chern insulators, including those realized in magnetically doped topological insulators [24].
Data availability— The data and code used to generate the figures are available at [25].
Acknowledgments — We thank Ulrike Nitzsche for her technical assistance. We acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through SFB 1170, Project-ID 258499086, and through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC2147, Project-ID 390858490). V.K. was funded by the European Union.